OpenStudy (anonymous):

3/5(30) find the product and reduce to lowest terms?

OpenStudy (angela210793):

3/(5*30) or (3/5) *30? or smth else? O.o

OpenStudy (anonymous):

Is it \[\frac{3}{5}50\]

OpenStudy (anonymous):

the way its written is 3/5(30) find product and reduce to lowest terms

OpenStudy (angela210793):

is 30 multiplying only 5 or all 3/5? i still don't get it

OpenStudy (anonymous):

no ishann it is (30) u have the 3/5 right

OpenStudy (anonymous):

\[\frac{3 \times 50}{5} or \frac{3}{50 \times 5}\]

OpenStudy (anonymous):

angel, either do i that's why I'm asking

OpenStudy (anonymous):

its not 50

OpenStudy (anonymous):

ok then just tell me the format

OpenStudy (anonymous):

is 30 above or below

OpenStudy (anonymous):

it is 3/5(30)....and then it says to find the product and then reduce to lowest terms

OpenStudy (angela210793):

maybe it is 90/5=18

OpenStudy (anonymous):

just like I wrote it

OpenStudy (anonymous):

angela, yes it says the answer is 18 hoever how do I get to that answer?

OpenStudy (anonymous):

however

OpenStudy (angela210793):

just multiply 3*30 and then divide by 5 3*30=90 90/5=18

OpenStudy (anonymous):

ohhh ok....been out of school 36 years, have forgotten sooo much

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I think when it's ambiguous where that (30) goes, you're supposed to read from left to right. In other words, you work out the \(\frac{3}{5}\) first, and then multiply by 30, which is numerically the same as \[\frac{3 \times 30}{5}\]

OpenStudy (angela210793):

:)

OpenStudy (anonymous):

Probable meant to be \[\frac{3}{5\times 30}=\frac{1}{5\times 10}=\frac{1}{50}\]

OpenStudy (angela210793):

he/she said the answer was 18...

OpenStudy (anonymous):

no stroodle, its 18, like angela and Dalvoron did

OpenStudy (anonymous):

a she here lol

OpenStudy (anonymous):

here's another.....5/18 divided by 15/16

OpenStudy (anonymous):

as you can see fractions are not by strong suit

OpenStudy (anonymous):

8/27?

OpenStudy (anonymous):

When dividing fractions, the quick and easy way is to flip the second fraction upside down, and chance the dividing sign to an equals sign \[\frac{5}{18} \div \frac{15}{16} = \frac{5}{18} \times \frac{16}{15}\]. Then just multiply the numerators together, and the denominators together to get a new fraction.

OpenStudy (anonymous):

\[=\frac{5}{18}\times\frac{16}{15}=\frac{1}{9}\times\frac{8}{3}=\frac{8}{27}\]

OpenStudy (anonymous):

ty all

OpenStudy (anonymous):

48/49*35/50*6/18=

OpenStudy (anonymous):

\[\frac{48}{49} \times \frac{35}{50} \times \frac{6}{18}\]